# Deriving pulsar pair-production multiplicities from pulsar wind nebulae using H.E.S.S. and LHAASO observations

### v1

@@ Abstract
Context. Pulsar Wind Nebulae (PWNe) dominate the galactic gamma-ray sky at very high energies, and are major contributors to the leptonic cosmic ray flux. However, whether or not pulsars also accelerate ions to comparable energies is not yet experimentally confirmed.
Aims. We aim to constrain the birth period and pair-production multiplicity for a set of pulsars. In doing so, we aim to constrain the proportion of ions in the pulsar magnetosphere and hence the proportion of ions that could enter the pulsar wind.
Methods. We estimate possible ranges of the value of the average pair production multiplicity for a sample of 26 pulsars in the Australia Telescope National Facility (ATNF) catalogue, which have also been observed by the High Energy Stereoscopic System (H.E.S.S.) telescopes. We then derive lower limits for the pulsar birth periods and average pair production multiplicities for a subset of these sources where the extent of the pulsar wind nebula and surrounding supernova shell have been measured in the radio. We also derive curves for the average pair production multiplicities as a function of birth period for sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO).
Results. We show that there is a potential for hadrons entering the pulsar wind for most of the H.E.S.S. and LHAASO sources we consider, dependent upon the efficiency of luminosity conversion into particles. We also present estimates of the pulsar birth period for six of these sources, which all fall into the range of $\simeq$10-50 ms.
@@ 1 Introduction
Pulsar wind nebulae (PWNe) occur when, following a supernova explosion, a pulsar’s ultra-relativistic wind flows into the surrounding medium producing a wind termination shock that can accelerate particles to relativistic energies (Gaensler & Slane, 2006; Kargaltsev et al., 2015). PWNe are the most numerous class of galactic very-high-energy gamma-ray emitters (representing around $\sim$40of known galactic sources), and have been observed across the electromagnetic spectrum for $\simeq$50 years (Wakely & Horan, 2008; H.E.S.S. Collaboration, 2018b). A key question regarding PWNe is whether the gamma-ray emission produced is solely a result of inverse Compton scattering of relativistic electrons on background fields, or if photopion production caused by protons which originated from the pulsar could contribute to the emission at the highest energies (Bednarek & Protheroe, 1997; Bednarek, 2003).
@@
A key metric which governs whether hadrons can likely escape the pulsar surface into the PWN is the average pair production multiplicity $\langle\kappa\rangle$. This describes the number of e+/e- pairs that escape the pulsar light cylinder per electron that escapes from the pulsar’s surface (as a result of cascades produced by gamma-ray Bremsstrahlung). This places some constraints on potential hadrons originating from the pulsar, as such particles do not multiply in cascades in the pulsar magnetosphere, and so can only make up a fraction of $1/\langle\kappa\rangle$ of the total particles at most (Kirk et al., 2007). Therefore for sources with a high pair-production multiplicity there exists the possibility of hadrons escaping into the wind. In this paper, we aim to determine the value of $\langle\kappa\rangle$ for a number of gamma-ray-emitting PWNe as observed by the High Energy Stereoscopic System (H.E.S.S.) gamma-ray telescopes (Aharonian, F. et al., 2006). This represents roughly a quarter of known PWN (Ferrand & Safi-Harb, 2012; Giacinti et al., 2020). We also make use of radio observations, as these can help to constrain the size of the PWN and surrounding Supernova Remnant (SNR) even once the associated gamma-ray emission has faded. Through this we can also place constraints on the pulsar birth period for a subset of these systems. We will also place constraints on the potential values of $\langle\kappa\rangle$ for a number of sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO) (Cao et al., 2022). This paper builds upon the previous work by de Jager (2007), who introduced the concept of deriving pair-production multiplicity constraints from gamma-ray data and comparing them to estimates of the pulsar birth period obtained from radio data, but we utilise a greater number of considered sources, updated observations and updated modelling.
@@
The remainder of this paper is organised as follows: in Section 2 we introduce the theoretical background to our models and discuss the input parameters to them, in Section 3 we present our main results. In Section 4 we discuss the implications for cosmic ray production, the effect of varying free parameters in the models and the prospects for future observations. Finally, in Section 5 we conclude.
@@ 2.1 Deriving $P_{0}$
Following the reasoning of van der Swaluw & Wu (2001), the initial spin period of a pulsar embedded in a PWN inside a larger SNR can be determined for systems aged $\sim 10^{3}-10^{4}$ yr using the relation
@@
where $E_{0}$ is the total mechanical energy of the Supernova Remnant (SNR) which we take to be $10^{51}\,\mathrm{erg}$, $I\approx 1.4\times 10^{45}\,\mathrm{g\,cm^{2}}$ is the moment of inertia of the neutron star (Mølnvik & Østgaard, 1985), $R_{SNR}$ and $R_{PWN}$ are the SNR and PWN radii as measured in the radio (as these mark the ‘true’ extent of the PWN (van der Swaluw & Wu, 2001)) and $P_{t}$ is the period of the pulsar at the present time. $\eta_{1}$ and $\eta_{3}$ are dimensionless scaling parameters that relate the radius of the PWN $R_{PWN}$ to the ratio of the SNR $R_{SNR}$
@@
where $E_{SD}$ is the total spin-down energy of the pulsar. $\eta_{1}$ effectively a proxy for the strength of synchrotron losses, neglecting these and setting $\eta_{1}=1$ and $\eta_{3}=1.02$ is suggested by van der Swaluw & Wu (2001), leading to a systematic underestimation of the initial spin rate which propagates through to our estimates of $\langle\kappa\rangle$. Equations 1 and 2 are only valid in the subsonic phase of the PWN’s evolution, after most of the energy in the pulsar wind has been deposited into the PWN, and follow the assumption that the PWN/SNR system is spherically symmetric (van der Swaluw & Wu, 2001). de Jager (2007) also considers the work of van der Swaluw & Wu (2001), although they ultimately adopt a fixed $P_{0}$ in their multiplicity analysis as a constraint of radio observations available at the time. We utilise a $P_{0}$ value calculated for each H.E.S.S. source for which we now have radio data in our analysis.
@@ 2.2 Determining $N_{el}$ and $\langle\kappa\rangle$
The number of Goldreich-Julian electrons at the current time $t$, which represents the total number of electrons that have been stripped from the polar caps of the pulsar, is given by the integral
$N_{GJ}=\int_{t=0}^{t=-\tau(P_{0})}\frac{[6c\dot{E}(t)]^{1/2}}{e}(-dt)\,$ (3)
(Goldreich & Julian, 1969). To calculate $\dot{E}$, we take values of $\dot{E}(t)$ today as calculated in Giacinti et al. (2020), and assume that the energy output of the pulsars evolve as
@@
with $\dot{E}_{t}$ the value of $\dot{E}$ at the current time, the constant $\tau_{0}$ set to $10^{3}\,$years and $\alpha=(n+1)/(n-1)$ assumed to be equal to 2 for a braking index $n=3$. $N_{GJ}$ is then integrated numerically. The average pair production multiplicity is then
$\langle\kappa\rangle=\frac{N_{el}}{2N_{GJ}}$ (5)
(de Jager, 2007).
@@
$N_{el}=E_{tot}\left(\frac{(2-\Gamma)E_{0}^{1-\Gamma}}{E_{2}^{2-\Gamma}-E_{1}^{2-\Gamma}}\right)\times\left(\frac{(E_{2}/E_{0})^{1-\Gamma}-(E_{1}/E_{0})^{1-\Gamma}}{1-\Gamma}\right)$ (6)
@@
where $\Gamma=2.2$, $E_{0}$ is assumed to be $0.1\,\mathrm{TeV}$, $E_{1}$ is the energy threshold of the gamma-ray observations (taken either from the H.E.S.S. galactic plane survey (H.E.S.S. Collaboration, 2018a) or (Aharonian et al., 2005) for G0.9+0.1/PSR J1747-2809) and $E_{2}$ is assumed to be $10\,\mathrm{TeV}$. This is as the radiative lifetime of electrons above this energy is likely to be less than the typical ages of pulsars in our sample, since this scales with electron energy $E_{e}$ as $\sim 10^{4}(B/10\,\mathrm{\mu G})^{-2}(E_{e}/10\,\mathrm{TeV})^{-1}\,\mathrm{yr}$ (Giacinti et al., 2020). The values for the total energy in electrons $E_{tot}$ are taken from the modelling in Giacinti et al. (2020). We then calculate the age $\tau$ as a function of $P_{t}$ the pulsar period today, the period derivative today $\dot{P}_{t}$, the initial period $P_{0}$ and the braking index $n$ such that
$\tau(P_{t},\dot{P}_{t},P_{0},n)=\left(1-\left(\frac{P_{0}}{P_{t}}\right)^{n-1}\right)\times\frac{P_{t}}{(n-1)\dot{P}_{t}}$ (7)
(Gaensler & Slane, 2006).
@@
Given the dependence of Equation 3 upon $\tau(P_{0})$ and therefore $P_{0}$ by Equation 7, we can find curves of $\langle\kappa\rangle$ as a function of $P_{0}$ (where $P_{0}$ is constrained to be shorter than the measured $P_{t}$ today). As these curves do not consider electrons producing emission outside the gamma-ray range, these curves serve as strict lower limits for the true pulsar multiplicity. This technique to determine these curves was first presented in (de Jager, 2007). As we can simultaneously find $P_{0}$ through the ratio of $R_{PWN}/R_{SNR}$ and Equation 1, for sources where $P_{t},R_{PWN},R_{SNR}$ are known and $N_{el}$ can be modelled, one can estimate a specific value for the lower bound of the pair production multiplicity. This is found by finding the intersection between the estimate for $P_{0}$ using Equation 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Equations 3, 6 and 7111Jupyter notebooks showing this analysis can be found at www.github.com/STSpencer/psrprops..
We also derive pair-production multiplicity curves for those sources observed by LHAASO with published results of modelling of their gamma-ray emission. Note that these publications make different assumptions about, for example, the associated radiation fields compared to the H.E.S.S. sources. For the Dragonfly Nebula, associated with PSR J2021+3651, we utilise a power law model with an exponential cutoff (PLEC), such that
[…]
following Woo et al. (2023). Here, we use a total energy of $3.9\times 10^{48}\,\mathrm{erg}$ to determine the proportionality constant numerically, a reference energy of $E_{1}=\mathrm{1\,TeV}$, a lower energy threshold of $E_{0}=\mathrm{25\,TeV}$, a cutoff of $E_{cut}=\mathrm{900\,TeV}$ and a maximum energy of $E_{2}=1400\,\mathrm{TeV}$ following Cao et al. (2023) and a spectral index of $\Gamma=\mathrm{1.4}$. For PSR J1841-0345 we also follow Albert et al. (2023) in using a power law spectrum with an exponential cutoff, with a proportionality constant of $7.1\times 10^{31}\,\mathrm{erg}$, a low-energy threshold of $E_{0}=\mathrm{10\,MeV}$, a reference energy of $E_{1}=\mathrm{7\,TeV}$, a cutoff energy of $E_{cut}=\mathrm{72\,TeV}$, a maximum energy of $E_{2}=\mathrm{740\,TeV}$ and a spectral index of $\Gamma=\mathrm{2.2}$. For PSR J1849-0001 we follow Amenomori et al. (2023) with a power-law model such that
$N_{el}\propto\left[\frac{1}{1-\Gamma}\left(\frac{E_{2}}{E_{0}}\right)^{1-\Gamma}-\frac{1}{1-\Gamma}\left(\frac{E_{1}}{E_{0}}\right)^{1-\Gamma}\right]$ (9)
with a low energy threshold of $E_{0}=\mathrm{0.5\,TeV}$, a reference energy of $E_{1}=\mathrm{10\,TeV}$, and high energy threshold of $E_{2}=\mathrm{100\,TeV}$ a spectral index of $\Gamma=2.5$ and a total energy of $2.8\times 10^{47}\,\mathrm{erg}$ from which the proportionality constant is also derived numerically. For PSR J1826-1334 we follow H. E. S. S. Collaboration et al. (2019) in using a broken power-law model with a two spectral parameters (BPL2) such that
[…]
with a low-energy threshold of $E_{0}=\mathrm{0.7\,TeV}$ (H.E.S.S. Collaboration, 2018a), a break energy of $E_{1}=\mathrm{0.9\,TeV}$, a maximum energy of $E_{2}=\mathrm{42\,TeV}$ a lower-energy spectral index of $\Gamma=1.4$ and a high-energy spectral index of $\Gamma_{2}=3.25$ with a total energy of $5.5\times 10^{48}\,\mathrm{erg}$ from which the proportionality constant is determined numerically. A summary of the input parameters to our modelling, along with the spectral models used for each source, are shown in Table 2. We obtained values for $R_{SNR}$ and $R_{PWN}$ from (Giacinti et al., 2020) and references therein.

### v3

@@ Abstract
Context. Pulsar wind nebulae (PWNe) dominate the galactic gamma-ray sky at very high energies and they are major contributors to the leptonic cosmic ray flux. However, the question of whether or not pulsars also accelerate ions to comparable energies has not yet been experimentally confirmed.
Aims. We aim to constrain the birth period and pair-production multiplicity for a set of pulsars. In doing so, we aim to constrain the proportion of ions in the pulsar magnetosphere and, hence, the proportion of ions that could enter the pulsar wind.
Methods. We estimated possible ranges of the value of the average pair production multiplicity for a sample of 26 pulsars in the Australia Telescope National Facility (ATNF) catalogue, which have also been observed by the High Energy Stereoscopic System (H.E.S.S.) telescopes. We then derived lower limits for the pulsar birth periods and average pair production multiplicities for a subset of these sources where the extent of the pulsar wind nebula and surrounding supernova shell have been measured in the radio. We also derived curves for the average pair production multiplicities as a function of birth period for sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO).
Results. We show that there is a potential for hadrons entering the pulsar wind for most of the H.E.S.S. and LHAASO sources we consider here, which is dependent upon the efficiency of luminosity conversion into particles. We also present estimates of the pulsar birth period for six of these sources, all falling into the range of $\simeq$10-50 ms.
@@ 1 Introduction
Pulsar wind nebulae (PWNe) follow a supernova explosion and the pulsar’s ultra-relativistic wind flows into the surrounding medium, producing a wind termination shock that can accelerate particles to relativistic energies (Gaensler & Slane, 2006; Kargaltsev et al., 2015). We know that PWNe are the most numerous class of galactic very-high-energy gamma-ray emitters (representing $\sim$40of known galactic sources) and have been observed across the electromagnetic spectrum for $\simeq$50 years (Wakely & Horan, 2008; H.E.S.S. Collaboration, 2018b). A key question regarding PWNe is whether the gamma-ray emission produced is solely a result of inverse Compton scattering of relativistic electrons on background fields or if pion production caused by protons originating from the pulsar could contribute to the emission at the highest energies (Bednarek & Protheroe, 1997; Bednarek, 2003).
@@
A key metric that governs whether hadrons can likely escape the pulsar surface into the PWN is the average pair production multiplicity $\langle\kappa\rangle$. This describes the number of e+/e- pairs that escape the pulsar light cylinder per electron that escapes from the pulsar’s surface (as a result of cascades produced by gamma-ray Bremsstrahlung). This places some constraints on potential hadrons originating from the pulsar, as such particles do not multiply in cascades in the pulsar magnetosphere; thus, they can only make up a fraction of $1/\langle\kappa\rangle$ of the total particles at most (Kirk et al., 2007). Therefore for sources with a low pair-production multiplicity there exists the possibility of hadrons escaping into the wind. In this paper, we aim to determine the value of $\langle\kappa\rangle$ for a number of gamma-ray-emitting PWNe as observed by the High Energy Stereoscopic System (H.E.S.S.) gamma-ray telescopes (Aharonian, F. et al., 2006). This represents roughly a quarter of known PWN (Ferrand & Safi-Harb, 2012; Giacinti et al., 2020). We also make use of radio observations, as these can help to constrain the size of the PWN and surrounding supernova remnant (SNR), even once the associated gamma-ray emission has faded. On this basis, we can also place constraints on the pulsar birth period for a subset of these systems. We also set constraints on the potential values of $\langle\kappa\rangle$ for a number of sources recently observed by the Large High Altitude Air Shower Observatory (LHAASO; Cao et al., 2022). This paper builds upon the previous work by de Jager (2007), who introduced the concept of deriving pair-production multiplicity constraints from gamma-ray data and comparing them to estimates of the pulsar birth period obtained from radio data; however, we utilise a greater number of considered sources, updated observations, and updated modelling.
@@
This paper is organised as follows: In Sect. 2, we introduce the theoretical background to our models and discuss the input parameters to them. In Sect. 3, we present our main results. In Sect. 4, we discuss the implications for cosmic ray production, the effect of varying free parameters in the models, and the prospects for future observations. Finally, in Sect. 5 we present our conclusions.
@@ 2.1 Deriving $P_{0}$
Following the reasoning of van der Swaluw & Wu (2001), the initial spin period of a pulsar embedded in a PWN within a larger SNR can be determined for systems aged $\sim 10^{3}-10^{4}$ yr using the relation
@@
where $E_{0}$ is the total mechanical energy of the SNR which we take to be $10^{51}\,\mathrm{erg}$, $I\approx 1.4\times 10^{45}\,\mathrm{g\,cm^{2}}$ is the moment of inertia of the neutron star (Mølnvik & Østgaard, 1985), $R_{SNR}$ and $R_{PWN}$ are the SNR and PWN radii as measured in the radio (as these mark the ‘true’ extent of the PWN; van der Swaluw & Wu 2001), and $P_{t}$ is the period of the pulsar at the present time. Then, $\eta_{1}$ and $\eta_{3}$ are dimensionless scaling parameters that relate the radius of the PWN $R_{PWN}$ to the ratio of the SNR $R_{SNR}$:
@@
Here, $E_{SD}$ is the total spin-down energy of the pulsar, while $\eta_{1}$ effectively a proxy for the strength of synchrotron losses, neglecting these and setting $\eta_{1}=1$ and $\eta_{3}=1.02$ is suggested by van der Swaluw & Wu (2001), leading to a systematic underestimation of the initial spin rate which propagates through to our estimates of $\langle\kappa\rangle$. Equations 1 and 2 are only valid in the subsonic phase of the PWN’s evolution, after most of the energy in the pulsar wind has been deposited into the PWN, and follow the assumption that the PWN+SNR system is spherically symmetric (van der Swaluw & Wu, 2001). de Jager (2007) also considered the work of van der Swaluw & Wu (2001), although they ultimately adopted a fixed $P_{0}$ in their multiplicity analysis as a constraint of radio observations available at the time. We utilised a $P_{0}$ value calculated for each H.E.S.S. source for which we now have radio data in our analysis.
@@ 2.2 Determining $N_{el}$ and $\langle\kappa\rangle$
The number of Goldreich-Julian electrons at the current time, $t$, which represents the total number of electrons that have been stripped from the polar caps of the pulsar, is given by the integral (Goldreich & Julian, 1969):
$N_{GJ}=\int_{t=0}^{t=-\tau(P_{0})}\frac{[6c\dot{E}(t)]^{1/2}}{e}(-dt)\,.$ (3)
To calculate $\dot{E}$, we took values of $\dot{E}(t)$ today, as calculated in Giacinti et al. (2020), and assumed that the energy output of the pulsars evolves as
@@
with $\dot{E}_{t}$ the value of $\dot{E}$ at the current time, the constant $\tau_{0}$ set to $10^{3}\,$years and $\alpha=(n+1)/(n-1)$ assumed to be equal to 2 for a braking index $n=3$. $N_{GJ}$ is then integrated numerically. The average pair production multiplicity is then (de Jager, 2007):
$\langle\kappa\rangle=\frac{N_{el}}{2N_{GJ}}.$ (5)
@@
$N_{el}=E_{tot}\left(\frac{(2-\Gamma)E_{0}^{1-\Gamma}}{E_{2}^{2-\Gamma}-E_{1}^{2-\Gamma}}\right)\times\left(\frac{(E_{2}/E_{0})^{1-\Gamma}-(E_{1}/E_{0})^{1-\Gamma}}{1-\Gamma}\right),$ (6)
@@
where $\Gamma=2.2$, $E_{0}$ is assumed to be $0.1\,\mathrm{TeV}$, $E_{1}$ is the energy threshold of the gamma-ray observations (taken either from H.E.S.S. Collaboration 2018a or Aharonian et al. 2005 for G0.9+0.1/PSR J1747-2809) and $E_{2}$ is assumed to be $10\,\mathrm{TeV}$. This is because the radiative lifetime of electrons above this energy is likely to be less than the typical ages of pulsars in our sample, since this scales with electron energy, $E_{e}$, as $\sim 10^{4}(B/10\,\mathrm{\mu G})^{-2}(E_{e}/10\,\mathrm{TeV})^{-1}\,\mathrm{yr}$ (Giacinti et al., 2020). The values for the total energy in electrons $E_{tot}$ are taken from the modelling in Giacinti et al. (2020). We can then calculate the age, $\tau$, as a function of $P_{t}$, the pulsar period today, the period derivative today, $\dot{P}_{t}$, the initial period, $P_{0}$, and the braking index, $n$, such that (Gaensler & Slane, 2006):
$\tau(P_{t},\dot{P}_{t},P_{0},n)=\left(1-\left(\frac{P_{0}}{P_{t}}\right)^{n-1}\right)\times\frac{P_{t}}{(n-1)\dot{P}_{t}}.$ (7)
@@
Given the dependence of Eq. 3 upon $\tau(P_{0})$ and therefore $P_{0}$ by Eq. 7, we can find curves of $\langle\kappa\rangle$ as a function of $P_{0}$ (where $P_{0}$ is constrained to be shorter than the measured $P_{t}$ today). As these curves do not consider electrons producing emission outside the gamma-ray range, these curves serve as strict lower limits for the true pulsar multiplicity. This technique to determine these curves was first presented in (de Jager, 2007). As we can simultaneously find $P_{0}$ through the ratio of $R_{PWN}/R_{SNR}$ and Eq. 1, for sources where $P_{t},R_{PWN},R_{SNR}$ are known and $N_{el}$ can be modelled, one can estimate a specific value for the lower bound of the pair production multiplicity. This is found by finding the intersection between the estimate for $P_{0}$ using Eq. 2, and the curves of $\langle\kappa\rangle(P_{0})$ as estimated using Eqs. 3, 6, and 7111Jupyter notebooks showing this analysis can be found at www.github.com/STSpencer/psrprops..
@@
We also derived pair-production multiplicity curves for those sources observed by LHAASO with published results of modelling of their gamma-ray emission. We note that these publications make different assumptions about, for example, the associated radiation fields compared to the H.E.S.S. sources. For the Dragonfly Nebula, associated with PSR J2021+3651, we utilised a power law model with an exponential cut-off (PLEC), namely,
[…]
following Woo et al. (2023). Here, we use a total energy of $3.9\times 10^{48}\,\mathrm{erg}$ to determine the proportionality constant numerically, a reference energy of $E_{1}=\mathrm{1\,TeV}$, a lower energy threshold of $E_{0}=\mathrm{25\,TeV}$, a cutoff of $E_{cut}=\mathrm{900\,TeV}$, and a maximum energy of $E_{2}=1400\,\mathrm{TeV}$ following Cao et al. (2023), along with a spectral index of $\Gamma=\mathrm{1.4}$. For PSR J1841-0345, we also followed Albert et al. (2023) by using a power-law spectrum with an exponential cutoff, with a proportionality constant of $7.1\times 10^{31}\,\mathrm{erg}$, a low-energy threshold of $E_{0}=\mathrm{10\,MeV}$, a reference energy of $E_{1}=\mathrm{7\,TeV}$, a cut-off energy of $E_{cut}=\mathrm{72\,TeV}$, a maximum energy of $E_{2}=\mathrm{740\,TeV,}$ and a spectral index of $\Gamma=\mathrm{2.2}$. For PSR J1849-0001, we followed Amenomori et al. (2023), with the power-law model expressed as
$N_{el}\propto\left[\frac{1}{1-\Gamma}\left(\frac{E_{2}}{E_{0}}\right)^{1-\Gamma}-\frac{1}{1-\Gamma}\left(\frac{E_{1}}{E_{0}}\right)^{1-\Gamma}\right].$ (9)
Here, we have a low energy threshold of $E_{0}=\mathrm{0.5\,TeV}$, a reference energy of $E_{1}=\mathrm{10\,TeV}$, a high-energy threshold of $E_{2}=\mathrm{100\,TeV,}$ a spectral index of $\Gamma=2.5,$ and a total energy of $2.8\times 10^{47}\,\mathrm{erg}$ (from which, the proportionality constant can be derived numerically). For PSR J1826-1334, we follow H. E. S. S. Collaboration et al. (2019) in using a broken power-law model with a two spectral parameters (BPL2), such that
[…]
Here, we have a low-energy threshold of $E_{0}=\mathrm{0.7\,TeV}$ (H.E.S.S. Collaboration, 2018a), a break energy of $E_{1}=\mathrm{0.9\,TeV}$, a maximum energy of $E_{2}=\mathrm{42\,TeV,}$ a lower-energy spectral index of $\Gamma=1.4,$ and a high-energy spectral index of $\Gamma_{2}=3.25,$ with a total energy of $5.5\times 10^{48}\,\mathrm{erg}$ (from which, the proportionality constant can be determined numerically). A summary of the input parameters to our modelling, along with the spectral models used for each source, are shown in Table 2. We obtained values for $R_{SNR}$ and $R_{PWN}$ from (Giacinti et al., 2020) and references therein.